OpenStudy (anonymous):

Can somebody please help me?? A.S.A.P 9. Which polynomial identity will prove that 72 = 64 + 8? A. Difference of Squares B. Difference of Cubes C. Sum of Cubes D. Square of a Binomial

ganeshie8 (ganeshie8):

hint : 64 = 4^3 8 = 2^3

OpenStudy (anonymous):

So sum of cubes? because the exponent is a 3 meaning cubed and then adding them is the sum.

OpenStudy (catlover5925):

@garciaash I found it!!

OpenStudy (anonymous):

@Catlover5925 THANK YOUUUU!!!!